OpenStudy (anonymous):

through: (4,-4), parallel to y=-x-4

OpenStudy (anonymous):

Your going to have to use point slope form here. Which is y-y1= m(x-x1)

OpenStudy (texaschic101):

not necessarily...you can use y = mx + b form...you just need to know the slope

OpenStudy (anonymous):

You have the point (4,-4) and m=-1 When you plug the points in you would have y -(-4) = -1(x - 4) two negatives make a positive so y +4 = -1(x - 4) Next you distribute the -1 to (x-4) and you would get y +4 = -x + 4 Then you subtract the 4 from the right side of the equation and subtract it from the left and you would get y=-x

OpenStudy (anonymous):

Well we do know the slope so we are just trying to find an equation that would be parallel to the original @texaschic101

OpenStudy (texaschic101):

we do...but does the asker know how to find the slope

OpenStudy (anonymous):

The slope is in front of X

OpenStudy (anonymous):

y = mx + b The m=slope and b=y - intercept

OpenStudy (texaschic101):

you are correct in that the slope is -1. And parallel lines will have the same slope y = mx + b slope(m) = -1 (4,-4)....x = 4 and y = -4 now we sub...we are looking for b, the y intercept, since we already know the slope is -1. y = mx + b -4 = -1(4) + b -4 = -4 + b 0 = b so basically your parallel equation is : y = -1x + 0...or just y = -x

OpenStudy (anonymous):

Thats what i just said...

OpenStudy (texaschic101):

yes..I did not say yours was wrong. All I said is that you can use the slope intercept formula....or you can use point slope formula (your suggestion). Either answer should come out the same.

OpenStudy (anonymous):

What ever lol this is easy peasy will you help me with the question i posted? @texaschic101